OpenStudy (anonymous):

FAN AND MEDAL*** eduelis29 A hybrid car can travel 40 miles per gallon. Approximately how many gallons of fuel will the car need to travel 20 km? [1 mile = 1.6 km] A)0.3 B)0.8 C)2.5 D)3.2

OpenStudy (anonymous):

A hybrid car can travel 40 miles per gallon. Approximately how many gallons of fuel will the car need to travel 20 km? [1 mile = 1.6 km] A)0.3 B)0.8 C)2.5 D)3.2

OpenStudy (anonymous):

ITS A?

OpenStudy (anonymous):

Yes.

OpenStudy (campbell_st):

so if 1 mile = 8/5 km then 1 km = 5/8 miles so to find the number of miles in 20 km 20 x 5/8 = then divide the answer by 40 hope it helps

OpenStudy (anonymous):